OpenStudy (anonymous):

Wondering if someone can talk me through this one: Calculate the derivative using the appropriate rule or combination of rules. f(x) = x(x^5 + 6)(x + 2)

OpenStudy (anonymous):

first of all multiply x with any of the next two terms..

OpenStudy (anonymous):

doesnt matter? ill do it with the second.

OpenStudy (anonymous):

ok... you can multiply x to the pranthesis and also another mul. and then start for derivativing or using chain rule. i use second one

OpenStudy (turingtest):

after distributing the x into one of the parentheses no chain rule, just product rule

OpenStudy (anonymous):

so i got: (x^2+2x)(x^5+6)

OpenStudy (anonymous):

f(x)=(x^6+6x) (x+2) f'(x) = (6x^5 + 6) (1) + (1)(x^6+6x) done :)

OpenStudy (anonymous):

my answer is: 5x^7+10x^5+10x^6+12x

OpenStudy (anonymous):

is that right?

OpenStudy (turingtest):

you didn't take the derivative yet, you just distributed

OpenStudy (anonymous):

explanation: (f(x)*g(x))' = f'(x)*g(x) + g'(x)*f(x)

OpenStudy (turingtest):

you can take the derivative now though :)

OpenStudy (turingtest):

there are many ways to do this problem...

OpenStudy (anonymous):

i did the product rule of: (x^2+2x)(x^5+6)

OpenStudy (turingtest):

ah okay, just that you didn't write out the middle step so I think you missed something

OpenStudy (anonymous):

@TuringTest look as this.. f(x)=(x^6+6x) (x+2) f'(x) = (6x^5 + 6) (1) + (1)(x^6+6x) done :) this is a reply you can see above... he/she hasn't explained anything...

OpenStudy (turingtest):

@Kashan you clearly did not get my message to you, check your inbox

OpenStudy (anonymous):

@yes I got it now... thanks...

OpenStudy (anonymous):

so now i just add out (6x^5 + 6) (1) + (1)(x^6+6x)

OpenStudy (anonymous):

so is my answer: x^6+6x^5+6x+6

OpenStudy (turingtest):

do you see where that came from?

OpenStudy (anonymous):

@bmelyk yes just solve it like simple algebraic expressions..

OpenStudy (anonymous):

yes i see how he came up with that equation using the product rule after facotring in the x.

OpenStudy (turingtest):

so yeah, now it's normal simplification

OpenStudy (anonymous):

can you factor that anymore?

OpenStudy (anonymous):

no its the answer...

OpenStudy (anonymous):

ok thxxxxx :)

OpenStudy (anonymous):

no problem.. :)